### v6

# Classification of compact manifolds with positive isotropic curvature

Abstract
We show the following result: Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature. Then $M$ is diffeomorphic to a spherical space form, or a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by a cocompact discrete subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$, or a connected sum of a finite number of such manifolds. This extends previous works of Brendle and Chen-Tang-Zhu, and improves a work of Huang. The proof uses Ricci flow with surgery on compact orbifolds, with the help of the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds.
Key words: Ricci flow with surgery, positive isotropic curvature, connected sum, ambient isotopy uniqueness, closed tubular neighborhood
MSC2020: 53E20, 53C21
1 Introduction
In our previous work [19], based on Brendle’s remarkable paper [7], we get a rough result on the classification of compact manifolds with positive isotropic curvature of dimension $n\geq 12$, from which a conjecture of Schoen and a conjecture of Gromov (and of Fraser), both in dimensions $n\geq 12$, follow. Now we can improve the main result in [19], and get a more precise topological classification of compact manifolds with positive isotropic curvature of dimension $n\geq 12$.
Theorem 1.1.
Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature. Then $M$ is diffeomorphic to a spherical space form, or a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by a cocompact discrete subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$, or a connected sum of a finite number of such manifolds.
(Of course, the cocompact discrete subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ referred to in the above theorem should act freely on $\mathbb{S}^{n-1}\times\mathbb{R}$; moreover, it is well known that $\text{Isom}(\mathbb{S}^{n-1}\times\mathbb{R})=\text{Isom}(\mathbb{S}^{n-1})\times\text{Isom}(\mathbb{R})$ for $n\geq 3$. By [24] the converse of Theorem 1.1 is also true.) This extends the main theorem in Chen-Tang-Zhu [10], which deals with the 4 dimensional case, and Theorem 1.4 in Brendle [7], which deals with the case that the $n$-manifolds contain no nontrivial incompressible $(n-1)$-dimensional space forms, where $n\geq 12$. The result is analogous to the classification of compact 3-manifolds with positive scalar curvature by Perelman et al.
As an application we partially verify a conjecture in [19].
Corollary 1.2.
Let $n\geq 12$ be even, $\Gamma$ be a finite subgroup of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, and $M$ be the total space of an orientable $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{S}^{1}$. Suppose that $M$ admits a metric with positive isotropic curvature. Then $M$ admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry.
(By $M$ admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry we mean that $M$ admits a Riemannian metric which is locally isometric to the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$; in this case $M$ is diffeomorphic to a quotient of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ by some isometries.)
The following result is a simple consequence of Theorem 1.1; compare the Main Theorem in [23].
Corollary 1.3.
Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature. Suppose that the fundamental group of $M$ is finite. Then $M$ is diffeomorphic to a spherical space form.
In particular, since any compact manifold which admits a non-collapsed, ancient Ricci flow must have finite fundamental group, cf. [1], it follows that any compact (connected) manifold of dimension $n\geq 12$ which admits a non-collapsed, ancient Ricci flow with positive isotropic curvature is diffeomorphic to a spherical space form. (Recall that by combining Perelman’s result that “there is no shrinking breathers other than gradient solitons” on closed manifolds (see p.9 of [32], cf. also Theorem 1.3 in [30]) with Theorem 1.4 in [31] one sees that any compact shrinking Ricci soliton of dimension $n\geq 12$ with positive isotropic curvature is isometric to a spherical space form.)
As in the proof of Corollary 2 in [10], by using Corollary 5 in [35] we also have the following
Corollary 1.4.
Let $M$ be a compact smooth manifold of dimension $n\geq 12$. If $M$ admits a metric with positive isotropic curvature, then $M$ also admits a locally conformally flat metric with positive scalar curvature.
(But the converse is not true; see [24].)
Recall that Micallef and Moore [23] showed that for a compact manifold $M$ of dimension $n\geq 4$ with positive isotropic curvature, we have the homotopy groups $\pi_{i}(M)=0$ for $2\leq i\leq[\frac{n}{2}]$; compare Theorem 4.6(ii) in Schoen and Yau [36]. The following corollary extends Micallef-Moore’s result when $n\geq 12$.
Corollary 1.5.
Let $M$ be a compact smooth manifold of dimension $n\geq 12$ which admits a metric with positive isotropic curvature. Then $\pi_{i}(M)=0$ for $22\leq i\leq n-$ 2.
In [24] Micallef-Wang showed that a compact manifold $M$ of even dimension $n=2m\geq 4$ with positive isotropic curvature has the second Betti number $b_{2}(M)=0$. The following corollary extends their result when $n\geq 12$ (for our result the dimension $n$ need not be even).
Corollary 1.6.
Let $M$ be a compact smooth manifold of dimension $n\geq 12$ which admits a metric with positive isotropic curvature. Then $b_{i}(M)=0$ for $22\leq i\leq n-$ 2.
In the process of proof of Theorem 1.1 we get a slightly more general result.
Theorem 1.7.
Let $(\mathcal{O},g_{0})$ be a compact connected orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. Then $\mathcal{O}$ is diffeomorphic to an orbifold connected sum of a finite number of spherical orbifolds with at most isolated singularities.
(By an obvious extension of Theorem 1.1 in [24] to the case of orbifold connected sum (cf. also [9]), certain version of the converse of Theorem 1.7 is also true.) This improves Theorem 1.3 in [19]. Compare Theorem 2.1 in [10]. We emphasize that here our notion of orbifold connected sums (see [19]) is stronger than that defined on p. 48 of [10]. (As in [10] we also allow self-connected sums; but of course our notion of self-connected sums is also stronger than that given in [10].)
As in [10], [7], and [19], the main tool in our proof is Ricci flow with surgery, aided by some topological arguments. In the 4-dimensional case the topological part of the proof of the main theorem in Chen-Tang-Zhu [10] used a result of McCullough [26], which says that any diffeomorphism of a 3-dimensional spherical space form is isotopic to an isometry. Similar statement for $(n-1)$-dimensional spherical space forms with $n\geq 5$ is not true in general.
One of our key observations is that for the topological part in the higher dimensional case, it suffices to use the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds. More precisely, using this uniqueness and a (rough) classification of spherical orbifolds with (nonempty) isolated singularities (due to [10] in the even ($\geq 4$) dimensional case and to [19] in the odd ($\geq 3$) dimensional case) one can prove fine properties of orbifold ancient $\kappa$-solutions and standard solutions when they contain caps with ends of type $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ with $\Gamma$ nontrivial, extending what were done in [7] and [19] for caps with ends of type $\mathbb{S}^{n-1}\times(-1,1)$, and extend/strengthen the definition of $\varepsilon$-caps in [7] and [19]. Then we can recognize the topology of components in the surgery process which are covered by canonical neighborhoods. Moreover, using the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds again, one can determine the topology of a connected sum of at most two spherical orbifolds with at most isolated singularities. (It is clear from our arguments in this paper that the use of McCullough’s result in the proof of the main theorem in Chen-Tang-Zhu [10] can be avoided; moreover, to get a similar classification of compact manifolds of dimension $5\leq n\leq 11$ with positive isotropic curvature, the only thing that remains to do is to show a suitable curvature pinching estimate, as Hamilton did in dimension 4 in [16] and Brendle did in dimensions $n\geq 12$ in [7].)
Recently, Zhengnan Chen [12] extended Brendle’s curvature pinching estimates to the case of $n\geq 9$. So combined with Chen’s estimates, our arguments imply that the above results also hold true for $9\leq n\leq 11$.
In Section 2 we define $\varepsilon$-caps, and describe the canonical neighborhood structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions. In Section 3 we construct an $(r,\delta)$-surgical solution to the Ricci flow starting with a compact, connected Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. As applications we prove Theorems 1.7 and 1.1, Corollaries 1.2, 1.3, 1.5 and 1.6, and derive a topological classification of compact Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. We will follow the notation and conventions in [19] in general.
2 $\varepsilon$-caps and ancient $\kappa$-solutions
We always use $\Gamma$ to denote a finite subgroup of $O(n)$ acting freely on $\mathbb{S}^{n-1}$.
To describe the structure of orbifold ancient $\kappa$-solutions and orbifold standard solutions we need notions on necks and caps.
Suppose $\sigma$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with at most isolated fixed points, let $\hat{\sigma}$ be the involution on the manifold $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ defined by $\hat{\sigma}(x,s)=(\sigma(x),-s)$ for $x\in\mathbb{S}^{n-1}/\Gamma$ and $s\in\mathbb{R}$, consider the quotient orbifold $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$, which has at most isolated singularities. We also denote this (Riemannian) orbifold by $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$; but beware that this notation has some ambiguity. Similarly we can define $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$. Note that we can consider $\Gamma$ and $\hat{\sigma}$ as isometries of $\mathbb{S}^{n}$ in a natural way, by lifting $\sigma$ to an isometry of $\mathbb{S}^{n-1}$ (which is always possible) and viewing $\mathbb{S}^{n}$ as a suspension of $\mathbb{S}^{n-1}$. We’ll use the same notation for these isometries of $\mathbb{S}^{n}$. The $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$ above is a smooth manifold if and only if $\sigma$ has no any fixed points in $\mathbb{S}^{n-1}/\Gamma$; if this is the case, it is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $B$ is a small, open metric ball centered at the point coming from the north and south poles of $\mathbb{S}^{n}$. Note that when $n$ is odd, $(\mathbb{S}^{n-1}/\Gamma\times\mathbb{R})//\langle\hat{\sigma}\rangle$ is a smooth manifold if and only if $\Gamma$ is trivial and $\langle\sigma\rangle$ acts on $\mathbb{S}^{n-1}$ antipodally; see the proof of Proposition 2.5 in [19]. Below when we write $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$, we always assume that the involution $\sigma$ has no any fixed points in $\mathbb{S}^{n-1}/\Gamma$.
If $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$ has nonempty isolated singularities, by Proposition 2.5 in [19], it must be diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$, where, as in [10], $\mathbb{S}^{n}//(x,\pm x^{\prime})$ denotes the quotient orbifold $\mathbb{S}^{n}//\langle\iota\rangle$ with $\iota$ the isometric involution on $\mathbb{S}^{n}$ given by $(x_{1},x_{2},\cdot\cdot\cdot,x_{n+1})\mapsto(x_{1},-x_{2},\cdot\cdot\cdot,-x_{n+1})$, which has two singular points, and $B$ is a small, open metric ball around a regular point in $\mathbb{S}^{n}//(x,\pm x^{\prime})$ such that the closure $\bar{B}$ of $B$ is disjoint from the two singular points and diffeomorphic to $D^{n}$. We’ll choose the regular point to be the image of the north pole under the natural projection $\mathbb{S}^{n}\rightarrow\mathbb{S}^{n}//(x,\pm x^{\prime})$.
Let $\mathcal{O}$ be a $n$-dimensional orbifold, $U$ an open subset, and $\Gamma$ as above. If $U$ is diffeomorphic to $\mathbb{S}^{n-1}//\Gamma\times\mathbb{R}$, we call it a topological neck. If $U$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$ or $\mathbb{S}^{n-1}//\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$, we call it a topological cap.
Let $\mathcal{O}$ be an orbifold, and $\mathcal{O}^{\prime}$ be a suborbifold of $\mathcal{O}$ in the sense of p. 108 in [21], so there is an immersion $f:\mathcal{O}^{\prime}\rightarrow\mathcal{O}$ such that $|f|$ maps $|\mathcal{O}^{\prime}|$ homeomorphically onto its image in $|\mathcal{O}|$. We identify $\mathcal{O}^{\prime}$ with its image in $\mathcal{O}$. Let $E(\xi)$ be the total space of an orbifold vector bundle $\xi$ over $\mathcal{O}^{\prime}$, and $\varphi:E(\xi)\rightarrow\mathcal{O}$ be a diffeomorphism onto an open neighborhood of $\mathcal{O}^{\prime}$ in $\mathcal{O}$, with the restriction of $\varphi$ to the zero-section of $\xi$ coinciding with the inclusion of $\mathcal{O}^{\prime}$ in $\mathcal{O}$. We call $\varphi$ an open tubular neighborhood of $\mathcal{O}^{\prime}$. Now we endow $\xi$ with a fiberwise inner product structure, so that it becomes a Euclidean orbifold vector bundle. Then we call the restriction of the above diffeomorphism $\varphi$ to the (unit) disk bundle $D(\xi)$ of $\xi$ a closed tubular neighborhood of $\mathcal{O}^{\prime}$. Compare Section 2 in Chapter VI of [6], Section 2 in Chapter III of [22], and Chapter 7 in [29].
Two open tubular neighborhoods $\varphi:E(\xi)\rightarrow\mathcal{O}$ and $\psi:E(\eta)\rightarrow\mathcal{O}$ of $\mathcal{O}^{\prime}$ is isotopic if there are a family of open tubular neighborhoods $\varphi_{t}:E(\xi)\rightarrow\mathcal{O}$ of $\mathcal{O}^{\prime}$ and a vector bundle isomorphism $\Lambda:E(\xi)\rightarrow E(\eta)$ such that $\varphi_{0}=\varphi$, $\varphi_{1}=\psi\circ\Lambda$, and the map $[0,1]\times E(\xi)\rightarrow\mathcal{O}$ taking $(t,v)\mapsto\varphi_{t}(v)$ is smooth. If $\xi$ and $\eta$ are endowed with fiberwise inner product structures, and the above $\Lambda$ can be taken to be an (orthogonal) isomorphism of Euclidean orbifold vector bundles, then we call the corresponding closed tubular neighborhoods of $\mathcal{O}^{\prime}$ isotopic. Compare Section 2 in Chapter VI of [6] and Chapter 7 in [29].
The following result on ambient isotopy uniqueness of closed tubular neighborhoods of a compact suborbifold in an orbifold should be known; cf. p.443 of [5]. It will be used extensively in this paper.
Proposition 2.1.
Let $\mathcal{O}^{\prime}$ be a compact suborbifold of an orbifold $\mathcal{O}$, and $\varphi:D(\xi)\rightarrow\mathcal{O}$ and $\psi:D(\eta)\rightarrow\mathcal{O}$ be two closed tubular neighborhoods of $\mathcal{O}^{\prime}$. Then there exists an isotopy $H_{t}$, $t\in[0,1]$, of $\mathcal{O}$ fixing $\mathcal{O}^{\prime}$ pointwisely, with compact support, and with $H_{0}=\text{id}$ and $H_{1}\circ\varphi=\psi\circ\Lambda$, where $\Lambda:D(\xi)\rightarrow D(\eta)$ is an orthogonal bundle isomorphism.
Proof.   First note that the two closed tubular neighborhoods $\varphi$ and $\psi$ are isotopic. This can be shown by using local models for the orbifold vector bundles $\xi$ and $\eta$; cf. the proof of Theorem 2.6 in Chapter VI in [6] and Theorem 7.4.6 in [29]. Alternatively, as pointed out to me by Prof. Bonahon, one can also show this by using normal exponential maps and the fact that the space of Riemannian metrics on the orbifold $\mathcal{O}$ is convex; cf. Section 6.1 of [2]. Then we can apply a version of Thom’s isotopy extension theorem for orbifolds. The (statement and) proof of the orbifold version of Thom’s isotopy extension theorem can be adapted from that of Theorem 9.5 in [8] and Theorem 7.3.3 in [29] (cf. also p. 179-181 of [25] and Theorem 2.4.2 in [37]). More precisely, the desired ambient isotopy in the orbifold version of Thom’s isotopy extension theorem comes from the projection of a suitable global flow on ×$\mathbb{R}\times\mathcal{O}$, the latter being generated by a vector field on ×$\mathbb{R}\times\mathcal{O}$. To construct this vector field, we first construct suitable vector fields in local models of ×$\mathbb{R}\times\mathcal{O}$ as in the proof of Theorem 9.5 in [8], Theorem 7.3.3 in [29] and Theorem 2.4.2 in [37], then get invariant vector fields in local models by using the method of averaging as in the proof of Theorem 3.1 in Chapter VI in [6], finally glue these local vector fields to a global vector field via a partition of unity. $\Box$
Let $D_{+}$ (resp. $D_{-}$) be the subset of $\mathbb{S}^{n}//\Gamma$ coming from points in $\mathbb{S}^{n}$ with $x_{n+1}\geq 0$ (resp. $x_{n+1}\leq 0$), and $p_{+}$ (resp. $p_{-}$) be the point in $D_{+}$ (resp. $D_{-}$) coming from the north (resp. south) pole of $\mathbb{S}^{n}$.
Lemma 2.2.
Let $\Gamma$ be as above, and $\mathcal{O}$ be a $n$-dimensional orbifold diffeomorphic to $\mathbb{S}^{n}//\Gamma$. Write $\mathcal{O}=\Omega_{1}\cup\Omega_{2}$, where $\Omega_{2}$ is diffeomorphic to $D_{-}$, and $\Omega_{1}\cap\Omega_{2}=\partial\Omega_{1}=\partial\Omega_{2}$. Assume that $h:\partial D_{-}\rightarrow\partial\Omega_{2}$ is a diffeomorphism which extends to a diffeomorphism $\tilde{h}:D_{-}\rightarrow\Omega_{2}$. Then there is a $\lambda\in\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$ such that $h\circ\lambda$ extends to a diffeomorphism $\tilde{h}^{\prime}:D_{+}\rightarrow\Omega_{1}$.
Proof.   Let $f:\mathbb{S}^{n}//\Gamma\rightarrow\mathcal{O}$ be a diffeomorphism. The case that $\Gamma$ is trivial follows from the Cerf-Palais disk theorem, and is well-known. Now we assume that $\Gamma$ is non-trivial. W.l.o.g. we may assume that $f(p_{-})\in\Omega_{2}$, since otherwise we can replace $f$ by $f$ precomposing with an isometric involution of $\mathbb{S}^{n}//\Gamma$ which exchanges $D_{+}$ and $D_{-}$. Then we may view both $f|_{D_{-}}:D_{-}\rightarrow f(D_{-})$ and $\tilde{h}:D_{-}\rightarrow\Omega_{2}$ as a closed tubular neighborhood of the singular point $f(p_{-})=\tilde{h}(p_{-})$. By the ambient isotopy uniqueness of closed tubular neighborhoods of compact suborbifolds (Proposition 2.1) there exist a $\Lambda\in\text{Isom}(D_{-})$ and a self diffeomorphism $F$ of $\mathcal{O}$ such that $F\circ f|_{D_{-}}=\tilde{h}\circ\Lambda$. Then $F\circ f|_{D_{+}}:D_{+}\rightarrow\Omega_{1}$ extends $h\circ\Lambda|_{\mathbb{S}^{n-1}/\Gamma}$, and we are done.
$\Box$
Remark. Note that the isometry group $\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$ may have more than two components when $\Gamma$ is nontrivial.
As before, let $B$ be a small, open metric ball in $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$ centered at the image of the north and south poles of $\mathbb{S}^{n}$ under the natural projection $\mathbb{S}^{n}\rightarrow\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$, and $\bar{B}$ be its closure.
Lemma 2.3.
Let $\Gamma$, $\sigma$, and $\hat{\sigma}$ be as before, and $\mathcal{O}$ be a $n$-dimensional orbifold which is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$. Write $\mathcal{O}=\Omega_{1}\cup\Omega_{2}$, where $\Omega_{2}$ is diffeomorphic to $\bar{B}$, and $\Omega_{1}\cap\Omega_{2}=\partial\Omega_{1}=\partial\Omega_{2}$. Assume that $h:\partial\bar{B}\rightarrow\partial\Omega_{2}$ is a diffeomorphism which extends to a diffeomorphism $\tilde{h}:\bar{B}\rightarrow\Omega_{2}$. Then there is a $\lambda\in\text{Isom}(\partial\bar{B})$ such that $h\circ\lambda$ extends to a diffeomorphism $\tilde{h}^{\prime}:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B\rightarrow\Omega_{1}$.
Proof.   The case that $\Gamma$ is trivial is treated in [19]. Now we assume that $\Gamma$ is nontrivial. Let $p$ be the (unique) singular point in $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$ coming from the north and south poles of $\mathbb{S}^{n}$. Let $f:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\rightarrow\mathcal{O}$ be a diffeomorphism. Then we may view both $f|_{\bar{B}}:\bar{B}\rightarrow f(\bar{B})$ and $\tilde{h}:\bar{B}\rightarrow\Omega_{2}$ as a closed tubular neighborhood of the singular point $f(p)=\tilde{h}(p)$. By Proposition 2.1 there exist a $\Lambda\in\text{Isom}(\bar{B})$ and a self diffeomorphism $F$ of $\mathcal{O}$ such that $F\circ f|_{\bar{B}}=\tilde{h}\circ\Lambda$. Then $F\circ f|_{\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B}:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B\rightarrow\Omega_{1}$ extends $h\circ\Lambda|_{\partial\bar{B}}$, and we are done. $\Box$
Now we start to analyze the structure of orbifold ancient $\kappa$-solutions which satisfy certain pinching assumption. (Recall from [7] and [19] that the orbifold ancient $\kappa$-solutions we consider are weakly PIC2.) Let $r_{1}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be the reflection w.r.t. the hyperplane $x_{1}=0$, we also use $r_{1}$ to denote the restriction of $r_{1}$ to $\mathbb{S}^{n-1}$. The following result is an improvement of Proposition 2.5 in [19]. (Recall that $C_{\text{PIC}}$, $C_{\text{PIC}1}$ and $C_{\text{PIC}2}$ are the sets of algebraic curvature tensors which have nonnegative isotropic curvature (weakly PIC, see item (i) of Definition 1.1 in [7]), weakly PIC1 (see item (ii) of Definition 1.1 in [7]), and weakly PIC2 (see item (iii) of Definition 1.1 in [7]) respectively.)
Proposition 2.4.
(cf. Corollary 6.7 in [7], Theorem 3.4 in [10], and Proposition 2.5 in [19])    Let $n\geq 4$, and $(\mathcal{O},g(t))$, $t\in(-\infty,T]$, be an orbifold ancient $\kappa$-solution of dimension $n$ with at most isolated singularities. Suppose that $(\mathcal{O},g(t))$ satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$ for some uniform constant $\theta>0$, and there is a spacetime point $(x_{0},t_{0})$ such that the curvature tensor at $(x_{0},t_{0})$ lies on the boundary of the cone $C_{\text{PIC}2}$. Then for each $t$, $(\mathcal{O},g(t))$ is isometric to a shrinking Ricci soliton $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ or $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}\mathbb{R}$. In particular, if $\mathcal{O}$ has no singularities and has exactly one end, it must be diffeomorphic to some $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, where $\sigma$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with no fixed points, and $\hat{\sigma}$ and $B$ are defined as before; if $\mathcal{O}$ has nonempty isolated singularities, it must be diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$ defined before. Furthermore, there exists a positive constant $\varepsilon_{4}=\varepsilon_{4}(n)\leq\varepsilon_{1}$, where $\varepsilon_{1}$ is the constant in Proposition 5.4 in [19], with the following property.
When $\mathcal{O}$ is diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$, any $2\varepsilon_{4}$-neck in $(\mathcal{O},g(t))$ must be diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$, and the central cross section $\Sigma$ of any $2\varepsilon_{4}$-neck in $\mathcal{O}$ bounds a compact domain $\Omega$; moreover, if $f:\mathbb{S}^{n-1}\rightarrow\Sigma$ is a $2\varepsilon_{4}$-homothety coming from the $2\varepsilon_{4}$-neck structure, and $f^{\prime}:\partial(\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B)\rightarrow\mathbb{S}^{n-1}$ is a homothety, either $f\circ f^{\prime}$ or $f\circ r_{1}\circ f^{\prime}$ extends to a diffeomorphism $F:\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B\rightarrow\Omega$.
When $\mathcal{O}$ is diffeomorphic to some manifold $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$ defined as above, any $2\varepsilon_{4}$-neck which is contained in the complement of the homothetically embedded $(\mathbb{S}^{n-1}/\Gamma)/\langle\sigma\rangle$ in $(\mathcal{O},g(t))$ must be diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(0,1)$, and the central cross section $\Sigma$ of such a $2\varepsilon_{4}$-neck bounds a compact domain $\Omega$ in $\mathcal{O}$; moreover, if $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ is a $2\varepsilon_{4}$-homothety coming from the $2\varepsilon_{4}$-neck structure, there exist some $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime}$ and a homothety $f^{\prime}:\partial\bar{B^{\prime}}\rightarrow\mathbb{S}^{n-1}/\Gamma$, such that $f\circ f^{\prime}$ extends to a diffeomorphism $F:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime}\rightarrow\Omega$.
Proof.   We only need to prove the last assertion when $\mathcal{O}$ is diffeomorphic to some manifold $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$ with $\Gamma$ nontrivial, as the remaining part is treated in Proposition 2.5 in [19]. In this case $\mathcal{O}$ is a smooth manifold with (only) one end, which is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(0,1)$, and $n$ must be even as $\Gamma$ is nontrivial (cf. the proof of Proposition 2.5 in [19]).
As in the proof of Proposition 2.5 in [19], we can write $\mathcal{O}=\mathbb{S}^{n-1}\times\mathbb{R}/\widetilde{\Gamma}$, where $\widetilde{\Gamma}$ is a subgroup of the isometry group of the round cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$ with $\widetilde{\Gamma}=\Gamma\cup\Gamma^{1}$, where the second component of $\Gamma$ (resp. of $\Gamma^{1}$) acts on $\mathbb{R}$ as the identity (resp. a reflection). Pick $\tilde{\sigma}\in\Gamma^{1}$. Then $\tilde{\sigma}^{2}\in\Gamma$, and $\tilde{\sigma}\Gamma=\Gamma^{1}$. So $\tilde{\sigma}$ induces an isometric involution on $\mathbb{S}^{n-1}/\Gamma$ (without any fixed points by our assumption on $\mathcal{O}$), which will be denoted by $\sigma$. The action of $\tilde{\Gamma}$ on $\mathbb{S}^{n-1}\times\mathbb{R}$ will leave exactly one cross section, say $\mathbb{S}^{n-1}\times\{0\}$, invariant. (Compare the beginning of the third paragraph on p. 53 in [10].) Let $\pi:\mathbb{S}^{n-1}\times\mathbb{R}\rightarrow\mathcal{O}=\mathbb{S}^{n-1}\times\mathbb{R}/\tilde{\Gamma}$ be the natural projection. Then $\pi(\mathbb{S}^{n-1}\times\{0\})$ is a homothetically embedded $(\mathbb{S}^{n-1}/\Gamma)/\langle\sigma\rangle$ in $(\mathcal{O},g(t))$. Fix any $t$ and write $g=g(t)$. From Proposition 5.4 in [19] we have the following
Claim. Assume that $0<\varepsilon_{4}=\varepsilon_{4}(n)\leq\varepsilon_{1}$. Any $2\varepsilon_{4}$-neck which is contained in the complement of $\pi(\mathbb{S}^{n-1}\times\{0\})$ in $(\mathcal{O},g)$ must be diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(0,1)$, and the central cross section $\Sigma$ of such a $2\varepsilon_{4}$-neck bounds a compact, connected, (smooth) submanifold in $\mathcal{O}$ which contains $\pi(\mathbb{S}^{n-1}\times\{0\})$.
Let $N$ be a $2\varepsilon_{4}$-neck contained in the complement of $\pi(\mathbb{S}^{n-1}\times\{0\})$ in $(\mathcal{O},g)$, by the Claim, it must be diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(0,1)$. Let $\psi:\mathbb{S}^{n-1}/\Gamma\times(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})\rightarrow N$ be a diffeomorphism giving the $2\varepsilon_{4}$-neck structure of $N$, and $\Sigma=\psi(\mathbb{S}^{n-1}/\Gamma\times\{0\})$. By the Claim again, $\Sigma$ bounds a compact domain, say $\Omega$, in $\mathcal{O}$ such that $\Omega$ contains $\pi(\mathbb{S}^{n-1}\times\{0\})$. Let $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ be a $2\varepsilon_{4}$-homothety coming from the $2\varepsilon_{4}$-neck structure, that is, $f$ is the same as $\psi|_{\mathbb{S}^{n-1}/\Gamma\times\{0\}}:\mathbb{S}^{n-1}/\Gamma\times\{0\}\rightarrow\Sigma$ after the canonical identification of $\mathbb{S}^{n-1}/\Gamma\times\{0\}$ with $\mathbb{S}^{n-1}/\Gamma$. As in the proofs of Proposition 6.17 in [7] and of Proposition 2.5 in [19], we will do metric surgery on $(\mathcal{O},g)$ along the $2\varepsilon_{4}$-neck $N$ to get a compact Riemannian orbifold $(\mathcal{O}^{\prime},g^{\prime})$ which is weakly PIC2 and strictly PIC if $\varepsilon_{4}=\varepsilon_{4}(n)$ is sufficiently small; moreover, there is a point in $\mathcal{O}^{\prime}$ where the Ricci curvature (even sectional curvature) is positive. We can write $\mathcal{O}^{\prime}=D_{1}\cup\Omega$, where $D_{1}$ is diffeomorphic to $D^{n}//\Gamma$, and $D_{1}\cap\Omega=\partial D_{1}=\partial\Omega=\Sigma$. By inspecting the surgery procedure we see that $f$ extends to a diffeomorphism from $D^{n}//\Gamma$ to $D_{1}$.
As in the proof of Proposition 2.5 in [19], by running the Ricci flow we see that $(\mathcal{O}^{\prime},g^{\prime})$ is diffeomorphic to a spherical orbifold with only one singular point, which has local group $\Gamma$, (and the complement of an open tubular neighborhood of the singular point is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B$,) so the spherical orbifold must be of the form $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle$ by using Lemma 5.2 in [10] and Corollary 2.4 in Chapter VI of [6], where $\sigma^{\prime}$ is an isometric involution of the spherical space form $\mathbb{S}^{n-1}/\Gamma$ with no fixed points, and the definition of $\hat{\sigma^{\prime}}$ is similar to that of $\hat{\sigma}$ above. Now we have a diffeomorphism $H:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle=\bar{B^{\prime}}\cup(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime})\rightarrow\mathcal{O}^{\prime}=D_{1}\cup\Omega$, where $B^{\prime}$ is a small, open metric ball centered at the (unique) singular point (coming from the north and south poles of $\mathbb{S}^{n}$). Let $f_{1}:\partial\bar{B^{\prime}}\rightarrow\mathbb{S}^{n-1}/\Gamma$ be a homothety. Note that $f\circ f_{1}$ extends to a diffeomorphism from $\bar{B^{\prime}}$ to $D_{1}$. Then the desired result follows from Lemma 2.3. $\Box$
Remark. It seems to be possible to choose $\sigma^{\prime}=\sigma$ in the above proposition with more efforts, maybe via torsion invariants (cf. [13] and [34]); but we don’t need this in this paper.
The following result extends Proposition 6.17 in [7] and Proposition 2.6 in [19].
Proposition 2.5.
(cf. Proposition 6.17 in [7] and Proposition 2.6 in [19]) Let $n\geq 4$ and $\varepsilon_{4}=\varepsilon_{4}(n)$ be as in Proposition 2.4. Let $(\mathcal{O},g)$ be a complete, noncompact orbifold diffeomorphic to $\mathbb{R}^{n}//\Gamma$ with strictly PIC and weakly PIC2 everywhere and strictly PIC2 at a point. Assume that $N$ is a $2\varepsilon_{4}$-neck in $\mathcal{O}$ diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(0,1)$, and $\Sigma$ is a central cross section of $N$. When $\Gamma$ is nontrivial we assume that $\Sigma$ bounds a compact domain $\Omega$ in $\mathcal{O}$ whose interior contains the unique singular point of $\mathcal{O}$. Then if $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ is a $2\varepsilon_{4}$-homothety coming from the $2\varepsilon_{4}$-neck structure of $N$, there exists a $\lambda\in\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$ such that $f\circ\lambda$ extends to a diffeomorphism $F:D^{n}//\Gamma\rightarrow\Omega$.
Proof.   The case that $\Gamma$ is trivial is treated in Proposition 2.6 in [19]. When $\Gamma$ is nontrivial, we can use a surgery argument as in the proof of Proposition 2.4, with the help of Lemma 2.2. $\Box$
The following definition of $\varepsilon$-caps extends/strengthens the corresponding definitions in [7] and [19], and is inspired by [7].
Definition. (cf. [7] and [19]) Let $\varepsilon_{0}=\varepsilon_{0}(n)$ be a small positive constant and $0<\varepsilon<\frac{1}{4}\varepsilon_{0}$. Let $(\mathcal{O},g)$ be a $n$-dimensional Riemannian orbifold with at most isolated singularities. Given a point $x_{0}\in\mathcal{O}$, an open subset $U$ of $\mathcal{O}$ is an $\varepsilon$-cap centered at $x_{0}$ if $U$ is a topological cap and $U\setminus V$ is an $\varepsilon$-neck, where $V$ is a compact domain in $U$ diffeomorphic to some $D^{n}//\Gamma$ or $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B$ or $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B$ with $\partial V$ a central cross section of an $\varepsilon$-neck contained in $U$ (i.e., $\partial V$ is the image of $\mathbb{S}^{n-1}/\Gamma\times\{0\}$ under the diffeomorphism associated to the $\varepsilon$-neck) and with $x_{0}\in\text{Int}\hskip 1.42262ptV$, (we will call $V$ a core of the cap $U$,) and in addition,
1. when $U$ is diffeomorphic to $\mathbb{R}^{n}//\Gamma$, if $\Sigma$ is a central cross section of an $\varepsilon_{0}$-neck in $U\setminus V$ with an $\varepsilon_{0}$-homothety $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ coming from the $\varepsilon_{0}$-neck structure, $\Sigma$ bounds a compact domain $\Omega$ in $U$, and $f\circ\lambda$ extends to a diffeomorphism $F:D^{n}//\Gamma\rightarrow\Omega$ for some isometry $\lambda$ of $\mathbb{S}^{n-1}/\Gamma$;
2. when $U$ is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, if $\Sigma$ is a central cross section of an $\varepsilon_{0}$-neck in $U\setminus V$ with an $\varepsilon_{0}$-homothety $f:\mathbb{S}^{n-1}/\Gamma\rightarrow\Sigma$ coming from the $\varepsilon_{0}$-neck structure, $\Sigma$ bounds a compact domain $\Omega$ in $U$, and there exist some $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime}$ and a homothety $f^{\prime}:\partial\bar{B^{\prime}}\rightarrow\mathbb{S}^{n-1}/\Gamma$, such that $f\circ f^{\prime}$ extends to a diffeomorphism $F:\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime}\rightarrow\Omega$;
3. when $U$ is diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus\bar{B}$, if $\Sigma$ is a central cross section of an $\varepsilon_{0}$-neck in $U\setminus V$ with an $\varepsilon_{0}$-homothety $f:\mathbb{S}^{n-1}\rightarrow\Sigma$ coming from the $\varepsilon_{0}$-neck structure, $\Sigma$ bounds a compact domain $\Omega$ in $U$, and $f\circ f^{\prime}$ extends to a diffeomorphism $F:\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B\rightarrow\Omega$ for some homothety $f^{\prime}:\partial\bar{B}\rightarrow\mathbb{S}^{n-1}$.
Proposition 2.6.
(cf. Theorem 6.18 in [7], Theorem 3.9 in [10], and Proposition 2.7 in [19])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$ and $C_{2}=C_{2}(n,\theta,\varepsilon)$, such that given any noncompact orbifold ancient $\kappa$-solution $(\mathcal{O},g(t))$ of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$ and is not locally isometric to an evolving shrinking round cylinder, for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$, which falls into one of the following categories:
(a) $U$ is an $\varepsilon$-neck centered at $(x_{0},t_{0})$,
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ and diffeomorphic to $\mathbb{R}^{n}//\Gamma$;
moreover, the scalar curvature in $U$ at time $t_{0}$ is (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$.
Proof.   The proof is almost the same as that of Theorem 2.7 in [19], with the help of Proposition 2.5. $\Box$
Now we have the following description of the canonical neighborhood property of orbifold ancient $\kappa$-solutions as in Proposition 2.9 in [19], but with the strengthened definition of $\varepsilon$-caps.
Proposition 2.7.
(cf. [33], Corollaries 6.20 and 6.22 in [7], Theorem 3.10 in [10], Proposition 2.9 in [19])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$, $C_{2}=C_{2}(n,\theta,\varepsilon)$ and $C_{3}=C_{3}(n,\theta)$ with the following property: Suppose $(\mathcal{O},g(t))$ is an orbifold ancient $\kappa$-solution of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$. Then either $(\mathcal{O},g(t))$ is compact and strictly PIC2 for any $t$ (hence diffeomorphic to a spherical orbifold), or for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ of $\mathcal{O}$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is a strong $\varepsilon$-neck centered at $(x_{0},t_{0})$,
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$;
moreover, the scalar curvature in $U$ at time $t_{0}$ satisfies the derivative estimates
$|\nabla R|<C_{3}R^{\frac{3}{2}}\hskip 22.76219pt\mbox{and}\hskip 22.76219pt|\frac{\partial R}{\partial t}|<C_{3}R^{2}.$
Proof.    One can argue as in the proof of Proposition 2.9 in [19], using Propositions 2.4 and 2.6. $\Box$
For orbifold standard solutions we have the following result.
Proposition 2.8.
(cf. [33], Corollary A.2 in [11], Corollary 9.3 in [7], and Proposition 2.10 in [19])   Given a small positive constant $\varepsilon$, there exist positive constants $C_{1}^{\prime}=C_{1}^{\prime}(n,\varepsilon)$, $C_{2}^{\prime}=C_{2}^{\prime}(n,\varepsilon)$ and $C_{3}^{\prime}=C_{3}^{\prime}(n)$ with the following property: For each space-time point $(x_{0},t_{0})$ on an orbifold standard solution $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(t))$ (with at most an isolated singularity), there is an open subset $U$ of $\mathbb{R}^{n}//\Gamma$ with $\overline{B(x_{0},t_{0},C_{1}^{\prime-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}^{\prime}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{\prime-1}R(x_{0},t_{0})$ and $C_{2}^{\prime}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is an $\varepsilon$-neck centered at $(x_{0},t_{0})$, and on $U\times[t_{0}-\min\{R(x_{0},t_{0})^{-1},t_{0}\},t_{0}]$ the solution $\hat{g}_{\Gamma}(t)$ is, after scaling with $R(x_{0},t_{0})$ and shifting $t_{0}$ to zero, $\varepsilon$-close to the corresponding subset of the evolving round cylinder $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ over the time interval $[-\min\{t_{0}R(x_{0},t_{0}),1\},0]$ with scalar curvature 1 at the time zero, and $U$ is disjoint from the surgery cap (in particular, $x_{0}\notin B(p_{\Gamma},0,\varepsilon^{-1})$) when $t_{0}\leq R(x_{0},t_{0})^{-1}$, or
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ (diffeomorphic to $\mathbb{R}^{n}//\Gamma$);
moreover, the scalar curvature in $U$ at time $t_{0}$ satisfies the derivative estimates
[…]
Proof.    The proof is almost the same as that of Proposition 2.10 in [19], with the help of Proposition 2.5. $\Box$
3 Existence of $(r,\delta)$-surgical solutions and applications
Now we can strengthen Proposition 3.2 in [19], and recognize the topology of a compact orbifold covered by $\varepsilon$-caps and/or $\varepsilon$-necks.
Proposition 3.1.
(cf. Proposition 3.2 in [19])   Let $0<\varepsilon_{0}\leq\varepsilon_{3}$, where the constant $\varepsilon_{3}$ is as in Proposition 5.7 in [19]. Fix $0<\varepsilon<\tilde{\varepsilon}_{1}(\varepsilon_{0})$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma 5.6 in [19]. Let $(\mathcal{O},g)$ be a closed, connected $n$-orbifold with at most isolated singularities. Suppose that each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck or an $\varepsilon$-cap. Then $\mathcal{O}$ is diffeomorphic to a spherical orbifold or a connected sum of at most two spherical orbifolds with at most isolated singularities. If we assume in addition that $\mathcal{O}$ is a manifold, then either $\mathcal{O}$ is diffeomorhic to a spherical space form, or $\mathcal{O}$ is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries.
Proof. Let $\mathcal{O}$ satisfy the assumption of our proposition. With our strengthened definition of $\varepsilon$-caps, we see that $\mathcal{O}$ is diffeomorphic to a spherical orbifold or a connected sum of at most two spherical orbifolds with at most isolated singularities; cf. the proof of Proposition 3.2 in [19]. Below we only indicate the cases where we can improve the arguments in [19].
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $\mathbb{R}^{n}//\Gamma$ ($|\Gamma|\geq 2$) but no $\varepsilon$-caps of the other types, by using the definition of $\varepsilon$-caps, Lemma 5.6 and Proposition 5.7 in [19], and Theorem 1.5 in Chapter 8 of [17], we see that it is diffeomorphic to $D^{n}//\Gamma\cup_{h}D^{n}//\Gamma$, where $h\in\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$, so it is diffeomorphic to a connected sum $\mathbb{S}^{n}//\Gamma\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points). In fact, it is diffeomorphic to $\mathbb{S}^{n}//\Gamma$; cf. Theorem 2.2 in Chapter 8 of [17] and Proposition 7.7 in [2].
Similarly, if $\mathcal{O}$ contains an $\varepsilon$-cap of type $\mathbb{R}^{n}//\Gamma$ ($|\Gamma|\geq 2$) and an $\varepsilon$-cap of type $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$, we see that it is diffeomorphic to some $(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime})\cup_{h}D^{n}//\Gamma$ (the notation is as in Section 2), where $h$ is a homothety between the boundaries. Hence it is diffeomorphic to a connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points). In fact, it is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle$; cf. Theorem 2.2 in Chapter 8 of [17] and Proposition 7.7 in [2].
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus\bar{B}$ ($|\Gamma|\geq 2$) but no caps of the other types, we see that it is a smooth manifold diffeomorphic to some $(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\setminus B_{1})\cup_{h}(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle\setminus B_{2})$, where $h$ is a homothety between the boundaries. Clearly it is diffeomorphic to a connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ (where the connected sum occurs at two singular points).
We claim that any manifold which is diffeomorphic to a connected sum
$\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ (here $\Gamma$ may be trivial; when $\Gamma$ is nontrivial we require that the connected sum occurs at two singular points) is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries. The reason is as follows. Let $f_{0}:[-n,n]\rightarrow[0,\infty)$ be a function satisfying the following conditions:
(1) $f_{0}$ is even;
(2) $f_{0}$ is smooth, concave and positive on $(-n,n)$;
(3) $f_{0}(s)=\sqrt{(n-1)(n-2)}$ for $s\in[-1,1]$; and
(4) there is some $s_{1}\in(1,n)$ such that $f_{0}(s)=\sqrt{(n-1)^{2}-(s-1)^{2}}$ for $s\in[s_{1},n]$.
Following Chapter 12 of [27], we argue the existence of the function $f_{0}$. Let $h_{0}(s)=\frac{s-1}{\sqrt{(n-1)^{2}-(s-1)^{2}}}$ for $s\in[0,n)$, and $\lambda:(-\infty,\infty)\rightarrow[0,1]$ be a non-decreasing smooth function which is identically 0 on $(-\infty,n-\frac{1}{2}+\nu]$ and identically 1 on $[n-\nu,\infty)$ for some positive number $\nu<\frac{1}{10}$. Note that
[…]
and
[…]
so there exists some $s_{0}\in(0,n-\frac{3}{2})$ such that
[…]
Setting
[…]
one can verify that $f_{0}$ is as desired.
Now let
[…]
be the revolutionary hypersurface in $\mathbb{R}^{n+1}$ obtained via rotating the graph of the function $f_{0}$; compare Chapter 12 in [27]. Then $\Sigma(f_{0})$ has an obvious $O(n)\times\mathbb{Z}_{2}$ symmetry, and $\Sigma(f_{0})$ is $O(n)\times\mathbb{Z}_{2}$-equivariantly diffeomorphic to the standard sphere $\mathbb{S}^{n}$. It follows that the spherical orbifold $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$ is diffeomorphic to the union of certain Riemannian manifold-with-boundary isometric to some $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$ (which is the quotient of $\mathbb{S}^{n-1}/\Gamma\times[-1,1]$ by the ${\mathbb{Z}_{2}}$-action generated by $\hat{\sigma}$ with $\hat{\sigma}(x,s)=(\sigma(x),-s)$ for $x\in\mathbb{S}^{n-1}/\Gamma$ and $s\in[-1,1]$) and a “standard” orbifold cap (which is the quotient of a “standard” (rotationally symmetric) cap (which has an end isometric to $\mathbb{S}^{n-1}\times[0,\mu]$ for some small $\mu>0$) by $\Gamma$) along their boundaries via an isometry between the boundaries. By combining this observation with the argument on p. 105 of [8] and Proposition 2.1, we see that any connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ (here $\Gamma$ may be trivial; when $\Gamma$ is nontrivial we require that the connected sum occurs at two singular points) is diffeomorphic to the union of two Riemannian manifolds-with-boundary along their boundaries via an isometry between the boundaries, where each of the Riemannian manifold-with-boundary is isometric to some $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$. (When $\Gamma$ is trivial, we use the Cerf-Palais disk theorem to replace Proposition 2.1.) Such a union has $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. (By the way, similarly, by using Proposition 2.1 and (the proof of) Theorem 2.2 in Chapter 8 of [17] (cf. also Proposition 7.7 in [2] and Theorem 10.2 in [20]) one can show that any connected sum $\mathbb{S}^{n}//\Gamma\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points when $|\Gamma|\geq 2$) is diffeomorphic to $\mathbb{S}^{n}//\Gamma$, and any connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points when $|\Gamma|\geq 2$) is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$.)
Thus the proof of Proposition 3.1 is finished. $\Box$
Now we adapt a definition from [4].
Definition (compare [4], [19]).    A piecewise $C^{1}$-smooth evolving compact Riemannian $n$-orbifold $\{(\mathcal{O}(t),g(t))\}_{t\in I}$ with at most isolated singularities is a surgical solution to the Ricci flow if it has the following properties.
i. The equation $\frac{\partial g}{\partial t}=-2\hskip 1.13809pt\text{Ric}$ is satisfied at all regular times;
ii. For each singular time $t_{0}$ there is a finite collection $\mathcal{S}$ of disjoint embedded $\mathbb{S}^{n-1}/\Gamma$’s in $\mathcal{O}(t_{0})$, and an Riemannian orbifold $\mathcal{O}^{\prime}$ such that
(a) $\mathcal{O}^{\prime}$ is obtained from $\mathcal{O}(t_{0})$ by removing a suitable open neighborhood of each element of $\mathcal{S}$ and gluing in a Riemannian orbifold diffeomorphic to $D^{n}//\Gamma$ along each boundary component thus produced which is diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$;
(b) $\mathcal{O}_{+}(t_{0})$ is a union of some connected components of $\mathcal{O}^{\prime}$ and $g_{+}(t_{0})=g(t_{0})$ on $\mathcal{O}_{+}(t_{0})\cap\mathcal{O}(t_{0})$;
(c) each component of $\mathcal{O}^{\prime}\setminus\mathcal{O}_{+}(t_{0})$ is diffeomorphic to a spherical orbifold, or a connected sum of at most two spherical orbifolds.
We fix $\varepsilon_{0}=\min\{\varepsilon_{3},\varepsilon_{4}\}$, where the constant $\varepsilon_{3}$ is as in Proposition 5.7 in [19], and the constant $\varepsilon_{4}$ is as in Proposition 2.4. Choose $0<\varepsilon<\frac{1}{4}\min\{\frac{1}{10^{3}n},\tilde{\varepsilon}_{1}(\varepsilon_{0})\}$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma 5.6 in [19].
Given a compact Riemannian orbifold $(\mathcal{O},g_{0})$ of dimension $n\geq 12$ with positive isotropic curvature and with at most isolated singularities, let $(\mathcal{O},g(t))$, $t\in[0,T)$, be the maximal solution to the (smooth) Ricci flow starting from $(\mathcal{O},g_{0})$. Set $\hat{T}=\frac{n}{2\inf_{x\in\mathcal{O}}R(x,0)}$. Then $T\leq\hat{T}$. By Theorem 1.2 in [7] there is a continuous family of closed, convex, $O(n)$-invariant sets $\{\mathcal{F}_{t}~{}|~{}t\in[0,\hat{T}]\}$ in the vector space $\mathcal{C}_{B}(\mathbb{R}^{n})$ of algebraic curvature tensors on $\mathbb{R}^{n}$ such that the curvature tensor of $(\mathcal{O},g_{0})$ lies in $\mathcal{F}_{0}$, the family $\{\mathcal{F}_{t}~{}|~{}t\in[0,\hat{T}]\}$ is invariant under the Hamilton ODE $\frac{d}{dt}\text{Rm}=Q(\text{Rm})$, and
$\mathcal{F}_{t}\subset\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}\}\cap\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}+f(R)\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}2}\}$
for any $t\in[0,\hat{T}]$, where $f:[0,\infty)\rightarrow[0,\infty)$ is an increasing concave function satisfying $\text{lim}_{s\rightarrow\infty}\frac{f(s)}{s}=0$, and $\theta$ is a positive number. As in [7], note that here $f$ and $\theta$ depend only on the initial data $(\mathcal{O},g_{0})$. By (a version of) Hamilton’s maximum principle the curvature tensor of $(\mathcal{O},g(t))$ lies in $\mathcal{F}_{t}$ for any $t\in[0,T)$. Let $\beta=\beta(\varepsilon)$ be the constant given in Lemma 5.8 in [19], and choose
$C=\max\{100\varepsilon^{-1},C_{1}(n,\theta,\varepsilon),C_{2}(n,\theta,\varepsilon),C_{3}(n,\theta),C_{1}^{\prime}(n,\beta\varepsilon),C_{2}^{\prime}(n,\beta\varepsilon),C_{3}^{\prime}(n)\},$
where the constants on the RHS are from Propositions 2.7 and 2.8.
We can also define the canonical neighborhood assumption $(CN)_{r}$ with $(4\varepsilon,4C)$-control and $(r,\delta)$-surgical solutions as in [33], [3], [4], and [19].
As in [19], with the help of Brendle’s curvature pinching estimates (see Theorem 1.2 and Corollary 1.3 in [7]), and Propositions 2.7, 2.8, and 3.1, we can show the following theorem.
Theorem 3.2.
(cf. Theorem 11.2 in [7], Theorem 4.6 in [10], and Theorem 4.11 in [19])    Let $(\mathcal{O},g_{0})$ be a compact Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. Let $\varepsilon$ and $C$ be chosen as above. Then we can find positive numbers $\hat{r}$ and $\hat{\delta}$ such that there exists an $(\hat{r},\hat{\delta})$-surgical solution (with $(4\varepsilon,4C)$-control) to the Ricci flow starting with $(\mathcal{O},g_{0})$, which becomes extinct in finite time.
Proof of Theorem 1.7. Theorem 1.7 follows from Theorem 3.2, Proposition 3.1, and Lemma 4.1 in [19]; cf. the proof of Theorem 1.3 in [19]. $\Box$
Proof of Theorem 1.1. Compare the proof of the Main Theorems in [10] and in [19]. Let $(M,g_{0})$ be a Riemannian manifold satisfying the assumptions of Theorem 1.1. From Theorem 1.7 we see that $M$ is diffeomorphic to a connected sum of a finite number of orbifolds, called components, where each component is diffeomorphic to either a spherical orbifold with at most isolated singularities or a connected sum of at most two spherical orbifolds with at most isolated singularities; furthermore, if a component is a manifold, by Proposition 3.1, it must be diffeomorhic to either a spherical space form, or a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries. We denote all components (whether being a manifold or not) by $\mathcal{O}_{1},\mathcal{O}_{2},\cdot\cdot\cdot,\mathcal{O}_{k}$.
First we undo the connected sum in the orbifold components which are diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{S}^{n}//(x,\pm x^{\prime})$ (where the connected sum occurs at two regular points) or $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{R}P^{n}$, and do orbifold connected sum operations among the orbifolds thus gotten and all other orbifold components to resolve the orbifold singularities of the components, which are created by the Ricci flow surgeries. (If a component is a smooth manifold, we don’t need to worry about it at this step.) By using the proof of Proposition 3.1 we see that what we get in this way are some manifolds, denoted by $Y_{j}$, where each of them is diffeomorphic to either a connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ or a self-connected sum of $\mathbb{S}^{n}//\Gamma$ at the two singular points. (Note that $\mathbb{S}^{n}//(x,\pm x^{\prime})$ is isometric to certain $\mathbb{S}^{n}//\Gamma$, where $\Gamma\cong\mathbb{Z}_{2}<O(n)$ acts on $\mathbb{S}^{n-1}$ antipodally.) By the last claim in the proof of Proposition 3.1 in the former case the resulting manifold admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. By an argument similar to that in the last claim in the proof of Proposition 3.1, we see that any self-connected sum of $\mathbb{S}^{n}//\Gamma$ (where $|\Gamma|\geq 2$) at the two singular points must also admit $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. (By the way, a similar argument with Proposition 2.1 replaced by Theorem 3.2 in Chapter 8 of [17] (see also Corollary 3.7 in Chapter III of [22] and Proposition 2.1 in [18]) shows that any self-connected sum of $\mathbb{S}^{n}$ also admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry; but this is well-known.)
Then we perform manifold connected sums to undo the Ricci flow surgeries which do not introduce orbifold singularities. So we do connected sum among $Y_{j}$, and a finite number of spherical space forms and a finite number of manifolds each diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries, which are components untouched in the previous step. We also do manifold connected sums which undo the decomposition of the orbifolds diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{S}^{n}//(x,\pm x^{\prime})$ or $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{R}P^{n}$ performed in the previous step. Thus we recover the original manifold $M$ as a desired manifold connected sum. $\Box$
Proof of Corollary 1.2. Let $M$ satisfy the assumptions of Corollary 1.2. Then $M$ has a connected sum decomposition as in the conclusion of Theorem 1.1. Note that $n$ is even and $M$ is orientable by assumption, the only orientable $n$-dimensional spherical space form is the sphere $\mathbb{S}^{n}$, and a manifold-with-boundary of the form $\mathbb{S}^{n-1}/\Gamma_{i}\times_{\mathbb{Z}_{2}}[-1,1]$ is non-orientable (recall that for even $n$ the action of any finite subgroup of $O(n)$ on $\mathbb{S}^{n-1}$ must be orientation-preserving if it is free), hence a manifold which is diffeomorphic to the union of two manifolds-with-boundary both of the form $\mathbb{S}^{n-1}/\Gamma_{i}\times_{\mathbb{Z}_{2}}[-1,1]$ along the boundaries is also non-orientable, so we can assume that in the connected sum decomposition of $M$ each summand is diffeomorphic to the total space of an orientable $\mathbb{S}^{n-1}/\Gamma_{j}$-bundle over $\mathbb{S}^{1}$ with structure group $\text{Isom}(\mathbb{S}^{n-1}/\Gamma_{j})$. By comparing the first Betti numbers we see that there is only one such summand, and we are done. $\Box$
Proof of Corollary 1.3. Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature and with $|\pi_{1}(M)|<\infty$. By the assumptions on $M$, we see that there is at most one nontrivial summand in the connected sum decomposition of $M$ given by the conclusion of Theorem 1.1, which must be diffeomorphic to a spherical space form. $\Box$
Proof of Corollary 1.5. Let $M$ satisfy the assumption of Corollary 1.5. By Corollary 1.2 in [19], which may also be viewed as a corollary of Theorem 1.1 here, there is a finite cover of $M$ which is diffeomorphic to $\mathbb{S}^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of a finite number of copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. From [28] we have an explicit construction of the universal cover, denoted by $\widetilde{M}$, of a connected sum of a finite number of copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. Observe that the image of a sphere $\mathbb{S}^{k}$ under a continuous map from $\mathbb{S}^{k}$ to $\widetilde{M}$ is compact for any $k\in\mathbb{N}$. By using induction, the Blakers-Massey theorem, and the long exact sequence for the relative homotopy groups of a pair, we see that $\pi_{i}(\widetilde{M})=0$ for $22\leq i\leq n-$ 2. Alternatively, once we know the universal cover $\widetilde{M}$, we can compute the homology groups $H_{i}(\widetilde{M})=0$ for $22\leq i\leq n-$ 2 (by using the Mayer-Vietoris sequence), and conclude $\pi_{i}(\widetilde{M})=0$ for $22\leq i\leq n-$ 2 by using the Hurewicz theorem. Now the corollary follows. $\Box$
Proof of Corollary 1.6. Let $M$ satisfy the assumption of Corollary 1.6. From the proof of Corollary 1.2 in [19], $M$ admits a finite sheeted regular cover, denoted by $\tilde{M}$, which falls into two cases: (1) $\tilde{M}$ is orientable, and is diffeomorphic to $\mathbb{S}^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$; (2) $\tilde{M}$ is non-orientable, and is diffeomorphic to $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ and $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. In the first case we have $H_{i}(\tilde{M})=0$ for $22\leq i\leq n-$ 2 by using the Mayer-Vietoris sequence, and
$H^{i}(M;\mathbb{R})\cong H^{i}(\tilde{M};\mathbb{R})^{G}=0,\hskip 22.76219pt2\leq i\leq n-2,$
(cf. for example, Example 2.6 in [15],) where $G$ is the deck transformation group of the covering map $\tilde{M}\rightarrow M$. In the second case, we take the orientation double cover, denoted by $\tilde{\tilde{M}}$, of $\tilde{M}$. Of course, an orientable double cover is a regular cover. Moreover we know that $\tilde{\tilde{M}}$ is diffeomorphic to $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. Now we can argue as in case (1) to conclude $H^{i}(\tilde{M};\mathbb{R})=0$ and $H^{i}(M;\mathbb{R})=0$ for $22\leq i\leq n-$ 2. $\Box$
Remark. In a preprint posted on the arXiv in January 2008, S. Gadgil and H. Seshadri pointed out that if the fundamental group of a compact manifold $M$ of dimension $n\geq 4$ with positive isotropic curvature is virtually free, then $b_{i}(M)=0$ for $22\leq i\leq n-$ 2.
As in [10], we also have a topological classification of compact Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature.
Theorem 3.3.
(cf. Corollary 5.3 in [10])    Let $(\mathcal{O},g_{0})$ be a compact connected Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. Then $\mathcal{O}$ is diffeomorphic to a connected sum of some spherical orbifolds with at most isolated singularities and some compact manifolds admitting $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry, where the connected sum occurs at regular points.
Proof. The result follows from Theorem 1.7 and the proof of Theorem 1.1 and Proposition 3.1. $\Box$
Acknowledgements. I would like to thank Prof. Francis Bonahon for answering my question on the paper [5] concerning tubular neighborhoods of suborbifolds. I would also like to thank Prof. Xu-an Zhao for helpful conversations. I was partially supported by NSFC (12271040) and Beijing Natural Science Foundation (Z190003).